wiki:UC20210202

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 02/02/2021

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Ben Williams (UW), Misty Bentz (GSU), Mukremin Kilic (OU), Joanne Hughes (Seattle U), Kal Kadlec (UW), Chip Kobulnicky (UWy), Sarah Tuttle (UW)


User feedback and comments from institutional representatives

  • Wyoming - nothing to report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle - nothing to report
  • Oklahoma - nothing to report
  • NMSU - nothing to report
  • NAPG - report sent by email; no updates other than having had recent bad luck with weather
  • JHU - nothing to report
  • Georgia State - nothing to report
  • FGCU - nothing to report
  • CU - nothing to report
  • BYU - no report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 1/05/2020 - 1/29/2021

0) Overview

APO has been operating at the 25% staffing limit since reopening on 12/2/2020. The State of New Mexico updated the tiered positions for every two weeks on Wednesday with the last update on 1/27.  Otero county remains in the Red Tier. The numbers spiked in January and are slowly going down; currently there is a 10.46% positivity rate and 34.9 cases per 100k people.  For comparison, Dona Ana county’s numbers from the same most recent state update were 9.46% positivity rate and 39.6 cases per 100k.

The data and numbers are available here: https://cv.nmhealth.org/public-health-orders-and-executive-orders/red-to-green/

We are still restricting all visitors to the site. Following the State of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors.

Weather for the month of January has been wintery, with cold temperatures and occasional snow. 

1) Telescope

The telescope had a failure of the altitude drive amplifier this past month.  The spare was put into operations and the failed one was sent to the manufacturer for evaluation and repair. 

2) Instruments

DIS: The scattered light in both the red and the blue cameras is stable and considered good to excellent.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light ratio (IOL) for red and blue continues to improve slowly with cooling weather.

NIC-FPS: powered on, cold and operational.

ARCTIC: powered on, cold, and operational. The issue with dome flats is being debugged and solutions are being worked on.  The design for the “fence” baffle on the mirror cover petal is being worked on; initial design ideas had to be modified because of interference issues with the telescope structure. The new redesigned rotation mechanism for the diffuser has arrived on site and installation will begin when staffing levels permit.  As a reminder, the ARCTIC shutter has issues when the temp approaches -10 degrees C. 

3) ARCSAT 

ARCSAT is operational and has been used for observations this past month.  We continue to investigate the issue between the two modes of operations for pointing. Off-axis guiding is available and recommended for exposures greater than 10 minutes. We are currently testing a remote controlled focuser for the guider. The Shelyak eShel preliminary commissioning took place at the beginning of November. The data from the commissioning work are still being analyzed and discussed. Further engineering is scheduled for Q1.


Additional telescope and instrument discussion

Nothing to add beyond what was included in Jamey's report above.


APO and the COVID-19 virus

All restrictions concerning visitors to the site and reduced staffing levels on site remain in place.


Q1 3.5m scheduling

There is still some open time available in Q1 (mostly A halves in March). Interested users should also check the availability of DD01 time in Q1. Users are reminded to follow the standard procedure of emailing requests for this time to Ben, Russet, Nancy, and their institutional scheduler.


Q1 ARCSAT scheduling

There is one open week available at the end of Q1.


Q2 3.5m scheduling

Ben will be sending out the institutional allocations soon; proposals will be due in late February.


KOSMOS Update

The UW TEG is in middle of tuning the thermal system in the KOSMOS dewar. This is the last step before installing the CCD into the dewar, which is currently in the shop because several copper connections have to be braised. The TEG is organizing plans for integration and testing in lab, e.g. by testing the newly manufactured slits, making sure the software system (with the TUI wrapper) works, and working on documentation. It is still difficult to provide an estimated ETA of the instrument on the mountain. Delivering an instrument when the mountain is still operating with significantly reduced staffing levels isn't useful because there would not be enough people on site to work on it. Six to eight weeks would be rushing, but it will likely be later than that.


ARCTIC Diffuser Rotation

There was an open window of time to work on the diffuser rotator in January, but most of that was lost due to illness. Bill is now working with Russet and Ben to figure out when they can next schedule some blocks of time with the diffuser off of ARCTIC to continue that work.


DAWG Report Feedback and Discussion

Last month the UC reps were asked to distribute the DAWG report to their department members and poll them for their highest priority items. This resulted in the following recommendations:

  • UW - strong interest quick look data pipelines and making sure that TUI has a longer lifetime
  • NMSU - porting TUI to python3, investigating options for a new (python-based?) version of Spextool for reducing TripleSpec data
  • Wyoming - porting TUI to python3, investigating hardware issues such as the fact that Apple's M1 chip is incompatible with some software
  • JHU - we should be moving away from IRAF for all instrument data reduction tools
  • CU - same as JHU, plus we should investigate cloud data storage such as AWS. Jamey reminded the group that we already have huge SDSS data transfers; additional data transfer to the cloud would tax the internet radio link (100 Mbit) on the mountain. We have investigated upgrading this link but it is currently cost prohibitive.
  • Seattle U - we should move away from IRAF ASAP but also need to assess the stability of python - potentially establish a data reduction advisory board to look into these issues?

Nancy will take these comments under advisement as she develops a strategy for moving forward.


ACTION ITEMS

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. Waiting on official quotes from vendor.
  • Develop a survey to get feedback on the pickup training session - Nancy. Status: OPEN. This is nearly done. Once it is completed, we will send it to those who were trained virtually in Q3 and Q4, as well as to the User Committee reps for distribution.

New action items from this meeting:

None.


All Other Business

None.


Next meeting

The next meeting will be on Tuesday March 2 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.